OpenStudy (anonymous):

Simplify: cos^2(x/4) - sin^2(x/4)

mathslover (mathslover):

\(\mathsf{\cos^2 A - \sin^2 A = \cos (2A) }\) Use this formula now. Here A = \(\cfrac{x}{4} \) .

mathslover (mathslover):

Can you do it now?

OpenStudy (anonymous):

Yeah thanks! I didn't think of the double angle formulas

mathslover (mathslover):

You're welcome :)